# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ VI Statistical metrics
Statistical metrics, like the means, variances, medians, and percentiles, are simple but significant components in the field of data analytics. Deriving statistical metrics has various real-world applications, and also acts as intermediate steps of many complex data analytics algorithms. The demand for computing the statistical metrics of client data with privacy preservation motivates numerous FA studies. In this section, we focus on three widely-used FA applications on statistical metrics: mean (Section VI-A), median/percentile (Section VI-B), and clustering center (Section VI-C).
@@ VI-A Means
As a strong and formal guarantee, LDP is applied in many FA-based mean computation solutions. LDP can be straightforwardly applied on mean computation by perturbing the client data $d_{i}$ before being uploaded. Based on the simple idea, researchers designed various solutions with their unique advantages. Cormode and Markov [21] propose an FA solution for the mean computation. In addition to satisfying the strong privacy criterion of LDP, their solution has a unique advantage, in that each client in their solution only needs to upload one bit to the server. Although it does not improve privacy preservation from a mathematical perspective, uploading only one bit is convincing for users with little mathematical backgrounds that the exposed information is minimized. In [22], an LDP solution is proposed to privately estimate the mean and histogram of distributed data. It also proposes a discretization technique to handle the continuous collection of user data (e.g., user data are collected daily). Harmony [23] provides an LDP solution for data analytics in mobile phone applications. It realizes mean estimation based on the Laplace mechanism, and frequency estimation based on the randomized response. PrivKV and its variations [24] extend the statistic tasks of mean estimation into the setting of data in the key-value structure. In that setting, each client holds a key-value pair; the data analyst expects to sort the keys of the clients and calculate the means of values with each same key. This paper then proposes an iterative structure to realize LDP with improved data utility. Honeycrisp [29] is a system designed to address large-scale count mean sketch tasks without a trusted server. A committee is randomly selected from users to generate homomorphic encryption keys. Client data is sent to the server using additively homomorphic encryption, which allows the server to calculate the total sum without accessing individual raw data. This sum is then decrypted by the committee. It utilizes the sparse vector theorem to manage the DP budget, enabling the committee to decrypt results and verify the aggregator’s calculations before deciding on their disclosure. In this case, the system also includes a summation tree method to ensure robustness and prevent the aggregator from cheating, where each node is the sum of its children under homomorphic encryption and they are verified by clients to ensure integrity.
@@
In the aggregation phase of FL, the FedAvg algorithm is computed on the client uploads. Essentially, the mean of gradient vectors from different clients is computed as the update on the global model. Many FL researchers try to improve the privacy and security of the aggregation phase, which falls into the category of FA-based mean computation.
@@
Secure aggregation-based approaches aim at improving the privacy preservation of uploading gradient vectors via cryptography tools. With secure aggregation, the server can only learn the aggregated result (mean of gradient vectors), but cannot obtain any knowledge about the uploads of any single client. Pairwise masking is the most lightweight and easy-to-implement technique to realize secure aggregation. In pairwise masking, each client communicates with some other clients, and a pairwise mask is generated for each pair of clients. The pairwise mask is added to the upload in one client in the pair, and its inverse is added to the other client in that pair. When an aggregator sums up the client uploads, the pairs of masks are eliminated, and the sum of the original data can be revealed. Fig. 6 provides an example of pairwise masking. In Fig. 6, three clients, holding local upload $X_{1},X_{2}$, and $X_{3}$ respectively, form three pairs and generate three pairwise masks. For example, mask $m_{1,2}$ is generated for clients 1 and 2, which is added to client 1’s upload, and inversely added to client 2’s upload. By summing up masked client uploads $U_{1},U_{2}$, and $U_{3}$, the sum of $X_{1},X_{2}$, and $X_{3}$ is revealed.
@@
Some other studies replace pairwise mask generation, which requires extensive client pair interactions, with non-pairwise masks. In non-pairwise mask generation, each client generates masks independently, and the server can eliminate the masks by learning the sum of the masks with a special protocol. In [119], a trusted third party is employed to assist with the masking, unmasking, and dropout handling. In LightSecAgg [120], the previous protocol is improved by removing the need for a trusted third party, based on a secure and dropout-resilience secret sharing scheme. In [121], the homomorphic pseudorandom generator (HPRG) is leveraged as a non-pairwise mask generation scheme to achieve lightweight, dropout-resilience secure aggregation.
DP-based approaches are also employed in privacy-preserving gradient aggregation. LDP is widely used in these solutions for its strong privacy guarantee. In [83], user-level DP is satisfied by applying Gaussian noise on the gradient vectors to enforce LDP. The proposed approach also provides a theoretical FL convergence guarantee, and trade-off between convergence speed and computation overhead. In [84], another LDP-enhanced aggregation approach is proposed by applying a Gaussian mechanism to the gradients. [84] also provides privacy-utility-communication trade-off, adjustable query sensitivity for LDP, and an accounting scheme of LDP budgets. LDP-FedSGD algorithm [84] considers privacy-preserving FL aggregation for IoT applications. It realizes LDP by perturbing the gradients with four LDP mechanisms. These mechanisms have different output structures and are optimal for different ranges of $\epsilon$ values. In HFL-DP [122], FL under the server-edge-client architecture is considered. The employed Gaussian mechanism provides an LDP guarantee under different levels on client uploads and edge uploads. Since the FL, and also the aggregation phases, take many rounds to complete, and the LDP budgets are cumulated by the composition theorem, naïvely applying LDP in each round could result in quite a low privacy budget for each round. Therefore, many of the aforementioned works [83, 84, 122] propose additional privacy-utility trade-offs or privacy budgeting schemes to tackle the issue.
On the other hand, since LDP typically introduces significant noise on the uploads, which harms the accuracy of mean computation results, some researchers investigate privacy-preserving gradient aggregation approaches satisfying the CDP criterion. In [87], the authors consider applying CDP on the aggregated gradient, based on a trusted aggregator (server). Such a CDP-enforced gradient can hide the participation of each participating client so that the privacy of each client is preserved. Gaussian noise is adopted to enforce CDP in [87]. In [88], another privacy-preserving aggregation scheme where similar settings of trusted aggregator, client-level privacy, and Gaussian mechanism, are considered. It is particularly designed for recurrent language models. The idea of CDP-enhanced aggregation is also employed in variations of FL like personalized FL [86], applying the Gaussian mechanism.
@@ VI-B Medians and percentiles
Median and percentile computation in the federated setting considers the clients each with a number of scaler data points $d_{i}$. $\mathcal{D}=\{d_{i}|i\}$ is the virtually aggregated data points from all clients. Consider a ranking function $\mathcal{R}(\mathcal{D},k)$ outputting the element in $\mathcal{D}$ exactly larger than $k$ portion of the elements in $\mathcal{D}$. The federated $k$-percentile computation lets the clients collaboratively derive the value of $\mathcal{R}(\mathcal{D},k)$, while preserving the privacy of data points $d_{i}$. The federated median computation is a special case of the federated $k$-percentile computation. The major challenge of federated median/percentile computation is that it requires different clients to compare their private data points to derive the ranking results.
As the private scaler computation is a quite classical problem setting in the field of MPC, some researchers design federated median/percentile solutions utilizing various cryptography tools. In [30], a private solution of median (and $k$-precentile) element over client local datasets is proposed. The solution is based on SMC and is realized by transforming the problem into a combinatorial circuit. In [31], a federated median/percentile computation scheme is proposed following the server-client architecture. The clients provide privatized input to the computation, and the server conducts the computation-intensive workloads to derive the results without observing the raw data. LDP is also employed as a formal privacy guarantee in federated median/percentile solutions. In [89], a median calculation scheme is proposed with both LDP and MPC. It can be extended to tasks like percentile computation. The LDP guarantee is realized by an exponential mechanism. The authors consider threat models of both semi-honest settings and adversarial settings. Another federated median/percentile computation solution [32] relies on a model-based optimization approach in a federated way, without a formal privacy guarantee. In [32] deriving the quantile is formulated as an optimization problem. That optimization problem is then collaboratively computed by the clients with ADMM. No formal privacy guarantee is provided.
@@ VI-C Clustering
The federated clustering task extends the clustering problem into the federated data setting. In that setting, the clients, each possessing a subset of data points, want to collaboratively derive the global clustering results, i.e., deriving what groups their data points belong to after the data points from different clients are virtually aggregated. The major challenge of federated clustering is that it needs to perform similarity measurements for data points from different clients while preserving data privacy.
Researchers have proposed federated clustering solutions following the FA paradigm. In these proposed algorithms, classical centralized clustering algorithms are adapted into the federated version, with the intermediate results serving as the FA insight. [17] proposes a federated clustering solution. It utilizes a hierarchical structure, where data points are first clustered locally within local client data, and local centroids are then clustered by the server. The authors derive an interesting finding, that in contrast to other federated computation solutions, data heterogeneity is beneficial to the analytics results. In [33], the authors consider kernel $k$-means, an effective clustering algorithm in capturing the nonlinear representation of the dataset using the nonlinear kernel functions, in the federated setting. To realize that, the clients first transform their data points into feature vectors by applying the kernel function. UIFCA [34] considers the unsupervised clustering tasks (the authors take unlabelled image clustering tasks as an example). It follows the idea of the IFCA algorithm, which trains a model for each cluster, and assigns a data sample to a cluster with the smallest model loss. To modify IFCA, which was originally designed for personalized FA, UIFCA replaces the models used to be generative models that capture the distribution of one cluster.

### v3

@@ VI Statistical Metrics
TABLE VII: Summary of statistical metrics solutions.
Reference Data analytics task Privacy Note
Cormode et al. [31] Mean computation LDP One-bit client upload
Ding et al. [32] Mean computation LDP Handling continuous data collection
Harmony [33] Mean computation LDP Laplace mechanism
PrivKV [34] Mean computation LDP Means of key-value structure data
Honeycrisp [37] Mean computation Cryptography&CDP Sparse vector theorem
FEVA [24] Mean computation Hiding raw data Federated Computation Builders architecture
SecAgg [38] Mean computation Cryptography Utilization of two kinds of masks
Bell et al. [39] Mean computation Cryptography Interacting with only a small random part of clients for each client
FastSecAgg [35] Mean computation Cryptography Multi-secret sharing scheme based on fast Fourier transform
Turbo [36] Mean computation Cryptography Circular communication topology
Zhao et al. [52] Mean computation Cryptography Trusted third party
LightSecAgg [53] Mean computation Cryptography a Secure and dropout-resilience secret sharing scheme
Liu et al. [54] Mean computation Cryptography Lightweight and dropout-resilience secure aggregation
Wei et al. [45] Gradient aggregation LDP Lightweight and dropout-resilience secure aggregation
Kim et al. [46] Gradient aggregation LDP Adjustable query sensitivity
HFL-DP [47] Gradient aggregation LDP Server-edge-client architecture
Geyer et al. [42] Gradient aggregation CDP Hiding the participation of each participating client
Mcmahan et al. [43] Gradient aggregation CDP Recurrent language models
Hu et al. [44] Gradient aggregation CDP CDP-enhanced aggregation in personalized FL
Aggarwal et al. [40] Medians and percentiles Cryptography Transforming the problem into a combinatorial circuit
Tueno et al. [41] Medians and percentiles Hiding raw data
Bohler et al. [51] Medians and percentiles Cryptohraphy& LDP Exponential Mechanism
[56] Medians and percentiles Hiding raw data Model-based optimization approach in a federated way
Dennis et al. [27] Clustering Hiding raw data Utilizing a hierarchical structure
Zhou et al. [58] Clustering Hiding raw data Applying kernel functions to transform data points into feature vectors
UIFCA [59] Clustering Hiding raw data Replacing the models used to be generative models that capture the distribution of one cluster
Lubana et al. [57] Clustering Hiding raw data Uploading both the cluster model parameters and the local centroids to the server
Servetnyk et al. [60] Clustering Hiding raw data Uploading a vector representing the number of samples falling into each bin of the grid
FedWalk [48] Graph metrics LDP FA version of a random walk
LF-GDPR [49] Graph metrics LDP Performing local computation by deriving the adjacency vector and the degree scaler each client
Liu et al. [50] Graph metrics LDP Protecting the information of vertex neighborhood within local graph data
In this section, we focus on FA solutions on the first class of data analytics tasks: statistical metrics. We consider four variations of statistical metrics problems: mean (Section VI-A), median/percentile (Section VI-B), clustering center (Section VI-C), and metrics in graph data (Section VI-D), and introduce these solutions in the four subsections. Statistical metrics, like the means, variances, medians, and percentiles, as well as metrics in graph data, are simple but significant components in the field of data analytics. Deriving statistical metrics has various real-world applications, and also acts as intermediate steps of many complex data analytics algorithms. The demand for computing the statistical metrics of client data with privacy preservation motivates numerous FA studies.
@@ VI-A Means
Many FA works are proposed for the sample analytics task of mean computation. In addition, some FA solutions tackle complex scenarios, but its mathematical model is essentially mean computation, like video analytics [24]. In the aggregation phase of FL, the FedAvg algorithm is computed on the client uploads. Essentially, the mean of gradient vectors from different clients is computed as the update on the global model. Many FL researchers try to improve the privacy and security of the aggregation phase, which falls into the category of FA-based mean computation [38, 39, 35].
Then, the FA mean computation solutions are introduced based on their privatization methodology, which is an effective taxonomy for FA mean computation tasks.
Local differential privacy method(s). As a strong and formal guarantee, LDP is applied in many FA-based mean computation solutions. LDP can be straightforwardly applied on mean computation by perturbing the client data $d_{i}$ before being uploaded. Based on the simple idea, researchers designed various solutions with their unique advantages. Cormode and Markov [31] propose an FA solution for the mean computation. In addition to satisfying the strong privacy criterion of LDP, their solution has a unique advantage, in that each client in their solution only needs to upload one bit to the server. Although it does not improve privacy preservation from a mathematical perspective, uploading only one bit is convincing for users with little mathematical backgrounds that the exposed information is minimized. In [32], an LDP solution is proposed to privately estimate the mean and histogram of distributed data. It also proposes a discretization technique to handle the continuous collection of user data (e.g., user data are collected daily). Harmony [33] provides an LDP solution for data analytics in mobile phone applications. It realizes mean estimation based on the Laplace mechanism, and frequency estimation based on the randomized response. PrivKV and its variations [34] extend the statistic tasks of mean estimation into the setting of data in the key-value structure. In that setting, each client holds a key-value pair; the data analyst expects to sort the keys of the clients and calculate the means of values with each same key. This paper then proposes an iterative structure to realize LDP with improved data utility.
LDP is also widely used for FA-empowered gradient aggregation in FL. In [45], user-level DP is satisfied by applying Gaussian noise on the gradient vectors to enforce LDP. The proposed approach also provides a theoretical FL convergence guarantee, and trade-off between convergence speed and computation overhead. In [46], another LDP-enhanced aggregation approach is proposed by applying a Gaussian mechanism to the gradients. [46] also provides privacy-utility-communication trade-off, adjustable query sensitivity for LDP, and an accounting scheme of LDP budgets. LDP-FedSGD algorithm [46] considers privacy-preserving FL aggregation for IoT applications. It realizes LDP by perturbing the gradients with four LDP mechanisms. These mechanisms have different output structures and are optimal for different ranges of $\epsilon$ values. In HFL-DP [47], FL under the server-edge-client architecture is considered. The employed Gaussian mechanism provides an LDP guarantee under different levels on client uploads and edge uploads. Since the FL, and also the aggregation phases, take many rounds to complete, and the LDP budgets are cumulated by the composition theorem, naïvely applying LDP in each round could result in quite a low privacy budget for each round. Therefore, many of the aforementioned works [45, 46, 47] propose additional privacy-utility trade-offs or privacy budgeting schemes to tackle the issue.
Central differential privacy method(s). Since LDP typically introduces significant noise on the uploads, which harms the accuracy of mean computation results, some researchers investigate privacy-preserving gradient aggregation approaches satisfying the CDP criterion. In [42], the authors consider applying CDP on the aggregated gradient, based on a trusted aggregator (server). Such a CDP-enforced gradient can hide the participation of each participating client so that the privacy of each client is preserved. Gaussian noise is adopted to enforce CDP in [42]. In [43], another privacy-preserving aggregation scheme where similar settings of trusted aggregator, client-level privacy, and Gaussian mechanism, are considered. It is particularly designed for recurrent language models. The idea of CDP-enhanced aggregation is also employed in variations of FL like personalized FL [44], applying the Gaussian mechanism.
@@
Cryptography method(s). Many researchers also apply cryptography tools to protect privacy in mean computation, which is termed secure aggregation in the scenarios of FL gradient aggregation. With secure aggregation, the server can only learn the aggregated result (mean of gradient vectors), but cannot obtain any knowledge about the uploads of any single client. Pairwise masking is the most lightweight and easy-to-implement technique to realize secure aggregation. In pairwise masking, each client communicates with some other clients, and a pairwise mask is generated for each pair of clients. The pairwise mask is added to the upload in one client in the pair, and its inverse is added to the other client in that pair. When an aggregator sums up the client uploads, the pairs of masks are eliminated, and the sum of the original data can be revealed. Fig. 8 provides an example of pairwise masking. In Fig. 8, three clients, holding local upload $X_{1},X_{2}$, and $X_{3}$ respectively, form three pairs and generate three pairwise masks. For example, mask $m_{1,2}$ is generated for clients 1 and 2, which is added to client 1’s upload, and inversely added to client 2’s upload. By summing up masked client uploads $U_{1},U_{2}$, and $U_{3}$, the sum of $X_{1},X_{2}$, and $X_{3}$ is revealed. To realize pairwise masking, the following works apply the novel hybrid model regarding its coordination model, so that masks can be exchanged via client-client interaction.
@@
Some other studies replace pairwise mask generation, which requires extensive client pair interactions, with non-pairwise masks. In non-pairwise mask generation, each client generates masks independently, and the server can eliminate the masks by learning the sum of the masks with a special protocol. In these non-pairwise mask schemes, the traditional server-client architecture is applied as the coordination model, which removes the extra requirement on clients to communicate with other clients. In [52], a trusted third party is employed to assist with the masking, unmasking, and dropout handling. In LightSecAgg [53], the previous protocol is improved by removing the need for a trusted third party, based on a secure and dropout-resilience secret sharing scheme. In [54], the homomorphic pseudorandom generator (HPRG) is leveraged as a non-pairwise mask generation scheme to achieve lightweight, dropout-resilience secure aggregation.
@@
The cryptography tools are also utilized in FA-based mean computation in general settings. Honeycrisp [37] is a system to address large-scale count mean sketch tasks without a trusted server. A committee is randomly selected from users to generate HE keys. Client data is sent to the server utilizing additive HE, which allows the server to calculate the sum without accessing individual raw data. This sum is then decrypted by the committee. It utilizes the sparse vector theorem to manage the DP budget, enabling the committee to decrypt results and verify the aggregator’s calculations before deciding on their disclosure. The system also includes a summation tree method to ensure robustness and prevent the aggregator from cheating, where each node is the sum of its children under HE and they are verified by clients to ensure integrity.
Hiding raw data method(s). In FEVA [24], an FA-based video analytics solution is proposed. It requires the clients to derive features from the video frames to preserve privacy following the hiding raw data methodology, and performs scaling and averaging procedures on the server side to derive the analytics result, based on the Federated Computation Builders architecture.
Lessons learned. Mean computation, which is the most fundamental task in statistics, gains diverse research interests, and many FA-based mean computation tasks are proposed. These applications either treat mean computation as a stand-alone task or resolve mean computation procedure in the FL aggregation framework (often termed secure aggregation), as the FL aggregation procedure is essentially a vector mean computation task. Various privatization techniques are applied in FA mean computation solutions, including LDP, CDP, and cryptography, under the trade-offs among accuracy, privacy, and computation/communication overheads. Extra issues in mean computation, including special data types, coordination model, and dropout resilience, are also considered by researchers.
@@ VI-B Medians and percentiles
Median and percentile computation in the federated setting considers the clients each with many scaler data points $d_{i}$. $\mathcal{D}=\{d_{i}|i\}$ is the virtually aggregated data points from all clients. Consider a ranking function $\mathcal{R}(\mathcal{D},k)$ outputting the element in $\mathcal{D}$ exactly larger than $k$ portion of the elements in $\mathcal{D}$. The federated $k$-percentile computation lets the clients collaboratively derive the value of $\mathcal{R}(\mathcal{D},k)$ while preserving the privacy of data points $d_{i}$. The federated median computation is a special case of the federated $k$-percentile computation. The major challenge of federated median/percentile computation is that it requires different clients to compare their private data points to derive the ranking results.
As private value ranking is a classical problem setting in the field of MPC, some researchers design federated median/percentile solutions utilizing various cryptography tools.
Cryptography method(s). In [40], a private solution of median (and $k$-precentile) element over client local datasets is proposed. The solution is based on MPC and is realized by transforming the problem into a combinatorial circuit. In [41], a federated median/percentile computation scheme is proposed following the server-client architecture. The clients provide privatized input to the computation, and the server conducts the computation-intensive workloads to derive the results without observing the raw data.
Local differential privacy + Cryptography method(s). In [51], a median calculation scheme is proposed with both LDP and MPC. It can be extended to tasks like percentile computation. The LDP guarantee is realized by an exponential mechanism. The authors consider threat models of both semi-honest settings and adversarial settings.
Hiding raw data method(s). Another federated median/percentile computation solution [56] relies on a model-based optimization approach in a federated way, without a formal privacy guarantee. In [56] deriving the quantile is formulated as an optimization problem. That optimization problem is then collaboratively computed by the clients with ADMM. No formal privacy guarantee is provided.
Lessons Learned. Computing medians and percentiles are different from other statistical metrics because the output is essentially an element of the input dataset. Therefore, the applications of DP mechanisms are limited, since adding random noise to the output is not desirable. As a result, DP applications on FA median/percentiles computation utilize the exponential mechanism, which adds noise to the output probabilities of the elements rather than the values of the elements. Cryptography tools that do not introduce randomness to the outputs also perform well in FA median/percentiles applications.
@@ VI-C Clustering
The federated clustering task extends the clustering problem into the federated data setting. In that setting, the clients, each possessing a subset of data points, want to collaboratively derive the global clustering results, i.e., deriving what groups their data points belong to after the data points from different clients are virtually aggregated. The major challenge of federated clustering is that it needs to perform similarity measurements for data points from different clients while preserving data privacy. Researchers have proposed federated clustering solutions following the FA paradigm. In these proposed algorithms, classical centralized clustering algorithms are adapted into the federated version, with the intermediate results serving as the FA insight. As for now, all existing FA-based clustering solutions only apply the idea of hiding raw data for privacy preservation. Designing FA-based clustering solutions with formal privacy preservation like DP and encryption is a valuable and promising research direction.
Hiding raw data method(s). [27] proposes a federated clustering solution. It utilizes a hierarchical structure, where data points are first clustered locally within local client data, and local centroids are then clustered by the server. The authors derive an interesting finding, that in contrast to other federated computation solutions, data heterogeneity is beneficial to the analytics results. In [58], the authors consider kernel $k$-means, an effective clustering algorithm in capturing the nonlinear representation of the dataset using the nonlinear kernel functions, in the federated setting. To realize that, the clients first transform their data points into feature vectors by applying the kernel function. UIFCA [59] considers the unsupervised clustering tasks (the authors take unlabelled image clustering tasks as an example). It follows the idea of the IFCA algorithm, which trains a model for each cluster, and assigns a data sample to a cluster with the smallest model loss. To modify IFCA, which was originally designed for personalized FA, UIFCA replaces the models used to be generative models that capture the distribution of one cluster.
@@
Lessons learned. FA clustering tasks derive the clustering results (e.g., centroids) from the data points held by different clients. Although clustering is an important statistics task, FA research on this topic is still somewhat primal. Existing solutions do not utilize advanced privatization mechanisms other than the fundamental idea of hiding raw data. These works convert the local data points into intermediate results of clustering (e.g., local clustering centroids) and transmit these results to the server for aggregation, which exposes remarkable privacy risks.
VI-D Graph metrics
The graph-structured data have yielded numerous promising graph analytics applications in the fields of social networks, multimedia, and AI4science. These applications require the utilization of novel graph metrics that are not present in conventional data structures. Leveraging FA to compute these graph metrics with privacy preservation can become a booster of graph analytics applications over edge data.
Local differential privacy method(s). In [49], a privacy-preserving graph metrics computation scheme named LF-GDPR is proposed. It supports various graph analytics tasks, including community detection and clustering coefficient estimation. Under the vertex-separated federated setting, each client performs local computation by deriving the adjacency vector and the degree scaler. Then, the insight is processed by LDP with careful privacy budget allocation. The derived metrics are aggregated on the server side to complete all kinds of graph analytics tasks utilizing dedicated aggregation algorithms proposed by the authors. [50] focus on the triangle counting problem in the graph-separated federated setting. The authors proposed an enhanced privacy notion named edge relationship LDP which can protect the information of vertex neighborhood within local graph data. The insight processed by edge relationship LDP therefore will not reveal sensitive information about the existence of edges.
Hiding raw data method(s). FedWalk [48] is an FA-based solution for graph embedding analytics. It derives expressive representations for the vertexes in a large graph. FedWalk considers the vertex-separated federated setting, where each client represents one vertex in the graph, and only possesses neighborhood information of the vertex. It performs the FA version of a random walk, a classical graph embedding learning algorithm.
Lessons learned. The graph data is a novel data structure with promising applications. The complex structure of a graph makes its privacy preservation quite different from conventional data, like scalar, vector, or discrete structures. Before applying particular privatization techniques, novel privacy criteria should be defined for the protection of edges, vertexes, or graphs, based on the particular data separation settings. After that, the corresponding FA graph analytics solutions are designed to optimize the performance of downstream graph analytics tasks.